OpenStudy (anonymous):

The equation T^2=A^3 shows the relationship between a planet’s orbital period, T, and the planet’s mean distance from the sun, A, in astronomical units, AU. If planet Y is twice the mean distance from the sun as planet X, by what factor is the orbital period increased? A. 2^1/3 B. 2^1/2 C. 2^2/3 D. 2^3/2

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

@Nnesha

OpenStudy (triciaal):

|sorry not 100% sure but I got C dw:1429767056107:dw|

OpenStudy (irishboy123):

\[\frac{T_x^2}{T_y^2}=\frac{A_x^3}{A_y^3}\] \[A_y = 2A_x\]